OpenStudy (anonymous):

Find an equation of the line through (2,4) and parallel to y=3x-7. Write the equation using function notation. f(x)=?

Nnesha (nnesha):

parallel lines have same slope

Nnesha (nnesha):

do you know what is slope of this equation y =3x - 7 ??

OpenStudy (anonymous):

3x

Nnesha (nnesha):

not 3x slope is just 3 :)

Nnesha (nnesha):

so like i said before parallel lines have same slope so what is slope of other equation

OpenStudy (anonymous):

you mean (2,4)?

OpenStudy (anonymous):

does y=3x-7= (3,-7)?

Nnesha (nnesha):

nope you have to find another equation which should parallel of this y =3x-7 so if slope of this equation is 3 then slope of other equation should 3 because lines are parallel

Nnesha (nnesha):

nope new equation should be y = 3x+b y = m x + b <--- slope intercept form

OpenStudy (anonymous):

ahh i forgot about mx+b

Nnesha (nnesha):

now use that order pair and replace x and y by their values (2,4) into y = 3x+b

OpenStudy (anonymous):

4=6+b

OpenStudy (anonymous):

do i divide all sides by 4?

Nnesha (nnesha):

looks good solve for b

Nnesha (nnesha):

nope 6 + b 6 is adding with b so do opposite

Nnesha (nnesha):

opposite of add is ??

OpenStudy (anonymous):

subtract

Nnesha (nnesha):

yes so subtract 6 both side

OpenStudy (anonymous):

-2=b

OpenStudy (anonymous):

or b=-2

Nnesha (nnesha):

yes now y = mx+b replace m and b by their values

OpenStudy (anonymous):

4=3(2)-2?

OpenStudy (anonymous):

f(x)=3x-2 ??

Nnesha (nnesha):

^^^this one

OpenStudy (anonymous):

so thats the correct one?

Nnesha (nnesha):

yes f(x)=3x-2 :)

OpenStudy (anonymous):

ok just making sure, THANK YOU, YOU JUST EARNED YOURSELF A NEW FAN :D

Nnesha (nnesha):

\(\huge\color{green}{{\rm my }~\rm pleasure !! :)}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\) :) and thanks :) my awesome fans!!!!1